Fermi National Laboratory
 

Accelerator Update for June 30, 2008 - July 02, 2008

 

Monday June 30

The day shift began with Operations monitoring the Tevatron (TeV) store 6255 (71.03E30) and the Recycler (RR) stash (369.3E10), with the Antiproton Source stacking (0), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the SciBooNE / MiniBooNE (BNB) experiments taking beam.

Operations terminated store 6255 at 9:06 AM.

Operations put the TeV into shot setup at 9:21 AM.

At 9:26 AM, Operations reported that a BLM at TeV sector B1 caused the TeV to abort with no beam in the machine. Everything reset.

At 9:43 AM, the TeV aborted for the first of two times in two minutes due to the B1 BLMs. Both times there was beam in the machine. Operations contacted an expert who suspected that the BLMs were in the wrong state. Operators rebooted “TEVCAL,” a front end that does the calculation that gives the BLM abort limit state settings. This fixed the problem.

At 10:50 AM, a Booster RF station (BRF9) tripped off. It reset.

At 11:04 AM, a Linac Klystron RF station (KRF2) tripped off on a LCW flow fault. It reset.

Water experts began cleaning the pond water strainers at MI-30 and MI-50 at 11:25 AM. The work took about 20 minutes.

Operations established store 6256 at 11:34 AM with an initial luminosity of 278.18E30.

A Pbar expert began adjusting the antiproton target at 12:23 PM.

Operations established beam to MTest at 12:24 PM.

At 12:45 PM, Operations reported the Pbar Debuncher kicker (D:EKIK) had tripped off and would not reset. Operations contacted an expert who investigated and reported that he found a bad output cable.

The Pbar kicker expert reported at 1:40 PM that he had D:EKIK back online.

Two MI experts began a multi- batch study period at 2:14 PM.

FESS personnel reported at 3:23 PM that they were beginning work on the air handler at AP-50. They were done 20 minutes later.

Operations reported at 4:03 PM that the Recycler had stashed antiprotons two times during the day shift.

At 4:26 PM, the BNB beam permit tripped off due to bad beam positions. It reset.

At 5:35 PM, Operations reported that BNB trips were occurring every five to ten minutes. Operators continued to investigate.

Operations disabled the MTest beam at 6:05 PM.

At 8:35 PM, BRF8 tripped off. It reset.

At 8:50 PM, Operations reported that a Pbar Debuncher RF station (DRF1-7) began tripping off on numerous “Driver Anode Current Not OK” indications. When Operations decided that they should contact an expert, the station resumed running normally.

At 10:24 PM, Operations reported that Booster Watchdog trips were inhibiting the Booster West Anode power supply. Operators investigated and then called in an expert.

Operations put Pbar into standby at 10:51 PM.

Booster resumed sending beam and Pbar resumed stacking at 11:37 PM.

A Booster expert reported at 11:59 PM that the WavTek generator used to generate the bunch rotation function had an internal fault. He switched to the backup generator.

Operations reported at midnight that the Recycler had stashed five times during the evening shift.

Tuesday July 1

The midnight shift began with Operations monitoring store 6256 (75.37E30) and the stash (285.3E10), with Pbar stacking (26.3E10), and with BNB taking beam.

At 12:50 AM, LRF4 tripped off. It reset.

Operations terminated store 6256 at 4:47 AM.

Operations put the TeV into shot setup at 4:58 AM.

At 5:34 AM, LRF3 tripped off for the first of two times in seven minutes. It reset.

Operations established store 6257 at 6:25 AM with an initial luminosity of 282.69E30.

Operations reported at 7:58 AM that the Recycler had stashed seven times during the midnight shift.

At 8:49 AM, Pbar experts reported that their Sweeper magnet “A” wasn’t working. Operations tried to switch to Sweeper magnet “B,” but it wasn’t working either.

A Pbar expert reported at 9:02 AM that he managed to bring Sweeper magnet “A” back on line. Sweeper “B” remained off, but only one Sweeper magnet was needed. Experts continued their investigation.

Pbar experts reported at 11:27 AM that they brought Sweeper magnet “B” online.

At 1:46 PM, a Booster vacuum pump (IPRF17) tripped off and took down BRF7 & 8. Everything reset.

A Booster expert reported at 3:35 PM that he had changed the timing module for a kicker (B:MP02) to a different control card.

Operations reported at 4:01 PM that the Recycler had stashed six times during the day shift.

At 4:39 PM, the Recycler 2-4 GHz vertical stochastic cooling system amp (R:V4A23) tripped off during a signal suppression measurement. Operators reset the system.

At 5:02 PM, Operations reported a tracking error for a Booster quadrupole (B:QS4).

Operators reported at 5:08 PM that they had search and secured the MI-65 enclosure.

Operations reported at 5:19 PM that NuMI experts had begun power supply testing.

At 6:27 PM, KRF3 tripped off. It reset.

At 7:50 PM, the BNB horn tripped off. It reset.

At 7:53 PM, a Booster vacuum pump (IPRF23) tripped off and brought down BRF13 & 14. Everything reset.

At 8 PM, KRF3 tripped off. It reset.

Operations terminated store 6257 at 8:06 PM.

Operations put the TeV into shot setup at 8:17 PM.

Operations brought the TeV sector D4 LCW pump online at 8:25 PM.

Operations established store 6258 at 9:54 PM with an initial luminosity of 283.72E30.

Operations reported at midnight that the Recycler had stashed four times during the evening shift.

Wednesday July 2

The midnight shift began with Operations monitoring store 6258 (183.39E30) and the stash (96.3E10), with Pbar stacking (25.0E10), and with BNB taking beam.

At 12:01 AM, BRF9 tripped off. It reset.

At 5:22 AM, the MI-50 power supply tripped off on a “Bucholz fast relay” indication. Operators investigated, but could find nothing wrong. The system would not reset, nor would the VCB close. Operations called in an expert.

At 6:35 AM, KRF6 tripped off. It reset.

The MI power supply expert had the MI-50 quadrupole power supply back on by 7:23 AM. Pbar resumed stacking. He said he could not find anything wrong either, but everything reset for him.

Operations reported at 7:57 AM that the Recycler had stashed four times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 07/02/2008    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory